OpenStudy (abb0t):

@blondie16 Algebra Question

OpenStudy (abb0t):

Simplify the expression: \[\frac{ (x^2-4)^\frac{ 1 }{ 2 }3-3x \frac{ 1 }{ 2 }(x^2-4)^{-\frac{ 1 }{ 2 }}2x }{ \left[ (x^2-4)^\frac{ 1 }{ 2 } \right]^2 }\]

OpenStudy (anonymous):

\[\frac{(x^2-4)^\frac{ 1 }{ 2 }(3)(x^2-4)^\frac{ 1 }{ 2 }-2x \sqrt{3} }{ (x-2)(x+2)(x^2-4)^\frac{ 1 }{ 2 } }\]

OpenStudy (anonymous):

there....^^^^^^^^ i doubt thats right though... LOL xDxD

OpenStudy (abb0t):

Not quite, but a good try. I can see how you got there.

OpenStudy (anonymous):

aweh mann :(

OpenStudy (anonymous):

lol maybe :P

OpenStudy (anonymous):

This should be a standard question for algebra exam

OpenStudy (anonymous):

answer is \[\frac{3}{\sqrt{x^2-4}}-\frac{3x^2}{(x^2-4)^\frac{3}{2}}\]

OpenStudy (anonymous):

i got \[\dfrac{-12}{(x^{2}-4)^\dfrac{3}{2}}\]

OpenStudy (anonymous):

\[\large{3x^2\over (x^2-4)^{3\over2}}+{3\over\sqrt{x^2-4}}\]

hartnn (hartnn):

\(\huge \color{Hotpink}{\checkmark }\)

OpenStudy (hba):

\[\huge\(3\sqrt{x^2-4}-\frac{ 3x^2 }{ x^2-4 })/(x^2-4)\] \[\huge\ 3(x^2-4)^{3/2}-3x^2(x^2-4)\] \[\huge\ 3x^3-24-3x^4+12x^2\] \[\huge\ -3x^4+3x^3+12x^2-24\] \[\huge\ -3(x^4+x^3-4x^2-8) < Answer\]

OpenStudy (jiteshmeghwal9):

One thing i wanna say everyone gave a diffrent a solution, who's right ?

OpenStudy (hba):

Well i guess i am right, i answered it right the last time. http://openstudy.com/users/abb0t#/updates/50dca993e4b0d6c1d542f834

hartnn (hartnn):

@hba what have you done ?? :P

hartnn (hartnn):

he solved it correctly in that link..... but not in this post :P

hartnn (hartnn):

\[\huge3\sqrt{x^2-4}-\frac{ 3x^2 }{\color{blue}{\sqrt{ x^2-4} }})/(x^2-4)\]

hartnn (hartnn):

yes, no.

OpenStudy (theviper):

Nice \(\Huge{\color{green}{L}\space\color{red}{a}\space\color{purple}{T}\color{pink}{e}\space\color{blue}{X}}\) @hba ;)

OpenStudy (hba):

Thanks :P

OpenStudy (abb0t):

I didn't think you guys would try it :P thanks guize. But I think raja has the correct answer.

OpenStudy (anonymous):

Woot Woot!